OpenStudy (anonymous):

Jeremy is making a 665 mile trip. He drives for 5 hours at 55 mph. What is the average rate of speed he must drive to complete the TOTAL trip in 11 hours? Assume that a steady speed is maintained for the rest of the trip.

OpenStudy (saifoo.khan):

(665 - (5*55)) / 6 = x

OpenStudy (saifoo.khan):

390 miles / 6 hours = 65 mile per hour

OpenStudy (anonymous):

5*55 = 275 665-275 = 390 390/6 = 65 65 MPH

OpenStudy (saifoo.khan):

i can see someone's cheating..

OpenStudy (saifoo.khan):

im NOT saying it jp-1203.

OpenStudy (anonymous):

you're helping me ;)

OpenStudy (saifoo.khan):

i did jp.. i'm saying it to SOMEONE else.

OpenStudy (saifoo.khan):

please dont take that on you! :)

OpenStudy (saifoo.khan):

Peter u here?